# Authorship Style Transfer with Policy Optimization

### v1

@@ Abstract
Authorship style transfer aims to rewrite a given text into a specified target while preserving the original meaning in the source. Existing approaches rely on the availability of a large number of target style exemplars for model training. However, these overlook cases where a limited number of target style examples are available. The development of parameter-efficient transfer learning techniques and policy optimization (PO) approaches suggest lightweight PO is a feasible approach to low-resource style transfer. In this work, we propose a simple two step tune-and-optimize technique for low-resource textual style transfer. We apply our technique to authorship transfer as well as a larger-data native language style task and in both cases find it outperforms state-of-the-art baseline models. [^1] Code, data, and models sufficient for a reproducibility study will be available at https://github.com/isi-nlp/ASTRAPOP/.
@@ 1 Introduction
In this work, we propose Authorship Style TRAnsfer with Policy OPtimization (Astrapop), a lightweight two-step PO training framework for authorship style transfer that only requires a single reward model, and which can be successfully transfer to an author’s style given as few as five examples. Unlike more complicated RL-based approaches like Hallinan et al. (2023), Astrapop is more computationally efficient and flexible, and works well with a variety of both RL-based and RL-free PO algorithms. We evaluate Astrapop on two authorship style transfer tasks, a low-resource individual authorship style transfer task and a medium-resource community authorship style transfer task. The evaluation results shows that Astrapop is more effectively able to leverage few-shot style transfer than ICL or SFT methods alone on the former task and also outperforms the state-of-the-art style transfer model with much less training time on the latter.
@@ 2.3.2 Paraphraser & Reference Model
We fine-tune off-the-shelf language models on the two generated supervised datasets to build the neutral paraphraser and the reference model. Specifically, for the neutral paraphraser, we simply fine-tune the model on dataset $\mathcal{N}$ to maximize
@@ 3 Experiments
We evaluate our approach on authorship style transfer tasks at two data resource levels, low-resource individual authorship style transfer and medium-resource community authorship style transfer. In this section, we first discuss the task specific details for each task, and then introduce the baseline models and implementation details.
@@ 3.1.4 Metrics
Content Preservation To measure content preservation, we use the SBERT [^3] We use the best-performing variant of SBERT, all-mpnet-base-v2. (Reimers and Gurevych, 2019) cosine similarity instead of the mutual implication score (MIS) (Babakov et al., 2022) in Patel et al. (2023) since MIS is trained on very short texts, but our test set contains much longer texts. The SBERT content preservation score is simply
@@ 3.2.1 Dataset
Our preliminary experiment shows that LLMs like GPT-3.5-turbo tend to drop information when paraphrasing long documents like the documents in the ETS dataset, so in this task, we process the documents at segment-level. Concretely, we split all documents into segments with up to 128 tokens, and the style label of the segments are the same as the original documents. We then sample 2,000 segments and 200 segments for each native language for training and validation, respectively, which results in a training set with 22,000 segments and a validation set with 2,200 segments. For evaluation, we transfer all documents in the test set into all native language styles except the source style to obtain transferred texts for all 110 native language pairs. This transfer is also done at segment-level. We segment all documents in the test set before transfer and regroup them back to documents afterward.
@@ 3.2.2 Transfer Model Formulation
Method Toward Away SBERT Joint Confusion
STRAP ${}^{\ddagger}$${}^{\dagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
STYLL 0.159 0.845 ${}^{\ddagger}$ ${}^{\ddagger}$ 0.59
SFT ${}^{\ddagger}$ ${}^{\ddagger}$ 0.754 ${}^{\ddagger}$ ${}^{\ddagger}$
Astrapop-PPO ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\dagger}$ ${}^{\dagger}$${}^{\dagger}$
[…]
Astrapop-CPO 0.165 ${}^{\ddagger}$ ${}^{\dagger}$${}^{\dagger}$ ${}^{\dagger}$
[…]
@@ 3.2.3 Reward Model
In the community-level transfer task, we use a classifier as the reward model $f_{reward}$ instead of the representation model since our preliminary experiment shows that the classifier works much better than the representation model on native language identification. Specifically, we train a RoBERTa-large (Liu et al., 2019b) classifier with 11 binary classification heads, corresponding to each native language on the ETS training set. Formally, given an input text x and a topic ($L1$) $s$, we denote the classifier output probability as $p_{s}(\textbf{x})$ and the classification decision as $C_{s}(\textbf{x})=\textbf{1}_{p_{s}(\textbf{x})>0.5}$. We then define the style similarity score as
@@ 3.2.4 Metrics
We use toward and away scores to indicate the the percentage increase in how many transferred documents are classified as being written by a target native language author and the percentage decrease in how many transferred documents are classified as being written by a source native language author. Formally, for each pair of source native language $s$ and target native language $t$, we define the toward score [^6] The toward score and the Toward objective/reward both measure to what extent the transferred texts reflect the target style, but they are defined differently. The toward score is defined to be more intuitive, while the Toward objective/reward is defined to be easier to calculate. Similar for the away score. as
@@ 3.4 Implementation Details
Zero-shot ${}^{\ddagger}$ ${}^{\dagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
STYLL ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
STRAP / SFT ${}^{\ddagger}$ ${}^{\ddagger}$ 0.917 ${}^{\ddagger}$ ${}^{\ddagger}$
STEER ${}^{\ddagger}$ 0.926 ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Astrapop-PPO ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$ ${}^{\ddagger}$
Astrapop-DPO ${}^{\dagger}$ ${}^{\ddagger}$${}^{\dagger}$${}^{\dagger}$${}^{\dagger}$
Astrapop-CPO 0.655 ${}^{\dagger}$ ${}^{\ddagger}$ 0.827 0.662
@@ 4 Results
In this section, we discuss and analyze the experimental results for both tasks. For conciseness, we only show the results with LLaMA-7B for all approaches and the results with the best reward combinations for Astrapop. Please see Appendix A for the full results and the ablation study. Due to the limited time and computational resources, we conduct all experiments in a single run, and perform statistical significance tests on the results. [^8] Please see § B.1 for details.
@@ 4.1 Individual Authorship Style Transfer
The automatic evaluation results on the individual authorship style transfer task are shown in Table 1. We only show the Astrapop models trained with the full reward function $R=T+A-(LP^{\alpha}-1)$ since this reward function yields the best Joint score for all three PO algorithms on this task. The joint score indicates that the overall performance of the three Astrapop models are superior to all baseline models, and the two models trained with the RL-free PO algorithms (i.e. DPO and CPO) perform similar to each other and both perform better than the RL-based PO algorithm (i.e. PPO).
Looking at the toward, away, and SBERT scores separately, we find that all PO algorithms can effectively improve the toward and the away scores, but at the cost of harming the SBERT score, since our reward function does not take semantic similarity into account, for efficiency and stability. Even so, the three Astrapop models still have decent SBERT scores that are higher than all baseline models except the SFT model since the KL-divergence penalty helps the models to preserve the capability to keep the semantic meaning of the input texts. One may notice that STYLL has much better away and confusion scores than all other models, but this is because the model sometimes copy some irrelevant content from the target exemplars which changes the meaning of the transferred texts, and this also explains why the SBERT score for STYLL is much lower than other models.
@@ 4.2 Community Authorship Style Transfer
The automatic evaluation results on the community authorship style transfer task is shown in Table 2. For this task, the best Astrapop models are trained with the reward function without the away reward $R=T-(LP^{\alpha}-1)$. The joint score indicates that DPO- and CPO-Astrapop have the best overall performance. They also has the best toward and confusion scores. PPO can also slightly improve the performance of the SFT model, but the improvement is much less than DPO and CPO. Similar to the previous task, PO training harms the SBERT score, but the magnitude of the loss is very small, and the result SBERT scores are still higher than all baseline models except STRAP/SFT.
We show a transfer example in Table 3 for a simple qualitative case study. It shows that Astrapop-CPO successfully captures a common typo, “alot” and three main characteristics of the target style: using all lowercase, using space before comma, and high typo rate, while no other model is able to capture any of these.
Target Sample
traviling is a very nice thing , it hepls you see new coultcurs and to meet new poeplo , there are alot of ways for traviling and i believe that the best way is to travil with a group led by a tour guide , [… more]
Source Text
The influnce of advertisements on the customers is worth commendable. The advertisers are projecting thier goods to the customers in a 'larger than real' manner.' [… more]
Zero-shot
Advertisements have a powerful influence on consumers, overstating the features and benefits of products to make them seem better than they actually are. [… more]
STYLL
Advertisements are used to promote products, making them appear attractive and useful. Through exaggeration, advertisers present their products as having exceptional features. [… more]
STRAP
The advertisements have a great effect on the customers and they should be praised. The advertisers make the products seem more better in the eyes of the customers rather than they really are. [… more]
STEER
Advertising has a positive impact on customers, as it promotes products in a way that exaggerates their qualities. Advertisers often portray their products as superior to reality. [… more]
Astrapop-CPO
because the advertisment have alot effec on the custumers , and make the products seem better than how they really are, [… more]
@@ 5 Related Work
Text Style Transfer Since parallel data is very rare for text style transfer, only a few works solve this task in a supervised manner (Zhu et al., 2010; Rao and Tetreault, 2018; Kim et al., 2022; Raheja et al., 2023). Constrained by the datasets, these works only focus on some specific sub-tasks such as text simplification and formality transfer. Therefore, to built more general style transfer models, recent works develop unsupervised methods that do not rely on parallel data. These works mainly fall in five categories, content-style representation disentanglement (Liu et al., 2019a; Dai et al., 2019), style-related phrase replacement (Madaan et al., 2020; Malmi et al., 2020; Reid and Zhong, 2021), reinforcement learning on direct objective (Gong et al., 2019; Liu et al., 2021b; Deng et al., 2022; Hallinan et al., 2023), pseudo-parallel data generation (Krishna et al., 2020; Riley et al., 2021), and LLM prompting (Reif et al., 2022; Suzgun et al., 2022; Patel et al., 2023).
The state-of-the-art authorship style transfer model, STYLL (Patel et al., 2023) transfers the input texts by prompting an LLM with the target style descriptors and few-shot pseudo-parallel transfer pairs generated by the same LLM, which combines the strength of pseudo-parallel generation and LLM prompting. Even so, as a prompting-based method, STYLL can be potentially enhanced by RL since RL has already been shown to be effective in improving the performance of prompting-based style transfer models (Deng et al., 2022), and the state-of-the-art general style transfer model, STEER (Hallinan et al., 2023) is also trained with RL. However, RL algorithms are shown to be unstable and hard to tune compare to the recently developed RL-free policy optimization algorithms such as DPO (Rafailov et al., 2023) and CPO (Xu et al., 2024).
Therefore, in this work, we choose the solve the authorship style transfer task with a PO-based training framework. Similar to STEER, we first generated pseudo-parallel data from the labeled non-parallel data and then train the model on the generated data, but our framework differs from STEER in three major ways: (1) we use a much simpler data generation strategy which only needs one paraphrase model and generates once for each instance in the non-parallel data, but STEER requires two extra models for each style as well as heavy overgeneration and filtering; (2) we only perform a single stage PO training instead of the two-stage offline-then-online RL training in STEER, and our reward function requires only one reward model instead of the three reward model in STEER; (3) we also use more stable and efficient RL-free PO algorithms instead of just the RL-based algorithm in STEER.
@@ 6 Conclusion
Although our approach shows strong performance on authorship style transfer, the performance on low-resource transfer is still much weaker than the performance on high-resource transfer. There are two possible reasons. First, we use small-scale datasets for both tasks due to the limited computational resources. It is sufficient to model the coarse-grained community authorship styles, but may be insufficient for the individual authorship styles. Therefore, if more computational resources are available, future work can investigate whether more training data can help improve the performance of the low-resource authorship transfer models. Second, our authorship information injection strategy may not be optimal. We use a popular exemplar-based approach to inject the authorship information in the low-resource transfer task, but there may be more efficient approaches such as using continuous vectors instead of discrete tokens. This is out of the scope of this work, but future work can explore more efficient information injection strategies for low-resource authorship style transfer.
Moreover, even though the two RL-free PO algorithms, DPO and CPO already show a much better performance than PPO, in this work, we only use them in an offline manner as in the original papers. However, the one can naturally enhance DPO and CPO with online data generated by the updated policy during training, which can potentially improve the performance of the models. Therefore, future work can focus on improving the training framework with online DPO or CPO training.
@@
Like other transfer learning LLMs, the quality of our model outputs highly depend on the quality of the underlying LLM and the training data. In this work, we use the original LLaMA-2-7B model instead of the chat version to ensure the flexibility for training, but it also has a higher risk of generating toxic texts. Also, the datasets we used contain unfiltered texts from the online forum Reddit and may also lead to unethical generation. Therefore, for real-world applications, we suggest carefully filtering the training data and also using a post-generation filter to avoid output unethical texts. As a PO-based training framework, one can also add some terms to the reward function to encourage the model to generate safe and ethical outputs.
Our model is intended for personal and authorized use such as building personal chatbot or authorship privatization, but we also recognize some potential harmful usage such as maliciously mimicking some individuals without authorization and intentionally generating texts in an offensive style. Therefore, we suggest keeping all personal data locally to prevent malicious mimicking. For text privatization, we suggest transferring to community-level authorship styles or styles mixed from multiple authors to prevent exposing information of individual authors. To maximally preclude any unintended use, we only permit the use of our approach on public datasets or with the explicit consent of the target authors.

### v2

@@ Abstract
Authorship style transfer aims to rewrite a given text into a specified target while preserving the original meaning in the source. Existing approaches rely on the availability of a large number of target style exemplars for model training. However, these overlook cases where a limited number of target style examples are available. The development of parameter-efficient transfer learning techniques and policy optimization (PO) approaches suggest lightweight PO is a feasible approach to low-resource style transfer. In this work, we propose a simple two-stage tune-and-optimize technique for low-resource textual style transfer. We apply our technique to authorship transfer as well as a larger-data native language style task and in both cases find it outperforms state-of-the-art baseline models. [^1] Code and models sufficient for a reproducibility study are available at https://github.com/isi-nlp/ASTRAPOP.
@@ 1 Introduction
In this work, we propose Authorship Style TRAnsfer with Policy OPtimization (Astrapop), a lightweight two-stage PO training framework for authorship style transfer, in which the first supervised fine-tuning stage prepares a reference model for the second stage, and the policy optimization in the second stage further improves the performance of the reference model by directly optimizing it on the authorship style transfer objective. Unlike more complicated RL-based approaches like Hallinan et al. (2023), Astrapop is more computationally efficient and flexible, and works well with a variety of both RL-based and RL-free PO algorithms. We evaluate Astrapop on two authorship style transfer tasks, a low-resource individual authorship style transfer task222with as few as five examples per author. and a medium-resource community authorship style transfer task. The evaluation results show that Astrapop is more effectively able to leverage few-shot style transfer than ICL or SFT methods alone on the former task and also outperforms the state-of-the-art style transfer model with much less training time on the latter.
@@ 2.3.2 Paraphraser & Reference Model
We fine-tune off-the-shelf language models on the two generated supervised datasets to build the neutral paraphraser [^3] Please see § B.5 for why we choose to train a local model instead of using GPT as the paraphraser. and the reference model. Specifically, for the neutral paraphraser, we simply fine-tune the model on dataset $\mathcal{N}$ to maximize
@@ 3 Experiments
We evaluate our approach on authorship style transfer tasks at two data resource levels, low-resource individual authorship style transfer and medium-resource community authorship style transfer. In this section, we first discuss the task specific details for each task and then introduce the baseline models and implementation details.
@@ 3.1.4 Metrics
Content Preservation To measure content preservation, we use the SBERT [^5] We use the best-performing variant of SBERT, all-mpnet-base-v2. (Reimers and Gurevych, 2019) cosine similarity instead of the mutual implication score (MIS) (Babakov et al., 2022) in Patel et al. (2023) since SBERT is the most commonly used semantic embedding model, and MIS is trained on very short texts, but our test set contains much longer texts. The SBERT content preservation score is simply
@@ 3.2.1 Dataset
Our preliminary experiment shows that LLMs like GPT-3.5-turbo tend to drop information when paraphrasing long documents like the documents in the ETS dataset, so in this task, we process the documents at segment level. Concretely, we split all documents into segments with up to 128 tokens, and the style labels of the segments are the same as the original documents. We then sample 2,000 segments and 200 segments for each native language for training and validation, respectively, which results in a training set with 22,000 segments and a validation set with 2,200 segments. For evaluation, we transfer all documents in the test set into all native language styles except the source style to obtain transferred texts for all 110 native language pairs. This transfer is also done at segment level. We segment all documents in the test set before transfer and regroup them back to documents afterward.
@@ 3.2.3 Reward Model
In the community-level transfer task, we use a classifier as the reward model $f_{reward}$ instead of the representation model since our preliminary experiment shows that the classifier works much better than the representation model on native language identification [^8] Please see § B.6 for detailed comparison.. Specifically, we train a RoBERTa-large (Liu et al., 2019b) classifier with 11 binary classification heads, corresponding to each native language on the ETS training set. Formally, given an input text x and a topic ($L1$) $s$, we denote the classifier output probability as $p_{s}(\textbf{x})$ and the classification decision as $C_{s}(\textbf{x})=\textbf{1}_{p_{s}(\textbf{x})>0.5}$. We then define the style similarity score as
@@ 3.2.4 Metrics
We use toward and away scores to indicate the percentage increase in how many transferred documents are classified as being written by a target native language author and the percentage decrease in how many transferred documents are classified as being written by a source native language author. Formally, for each pair of source native language $s$ and target native language $t$, we define the toward score [^9] The toward score and the Toward objective/reward both measure to what extent the transferred texts reflect the target style, but they are defined differently. The toward score is defined to be more intuitive, while the Toward objective/reward is defined to be easier to calculate. Similar for the away score. as
@@ 3.4 Implementation Details
[…]
STYLL 0.159 0.845 0.529 0.440 0.59
SFT 0.137 0.707 0.754 0.484 0.32
[…]
@@ 4 Results
In this section, we discuss and analyze the experimental results for both tasks. Due to the limited time and computational resources, we conduct all experiments in a single run and perform statistical significance tests on the results. [^11] Please see § B.1 for details. For conciseness, we only show the automatic evaluation results with LLaMA-7B for all approaches and the results with the best reward combinations for Astrapop. Please see § A.1 for the full automatic evaluation results and the ablation study. We also conduct human evaluation and a case study for the community authorship transfer experiments. Please see § A.4 and § A.5 for details.
@@ 4.1 Individual Authorship Style Transfer
The automatic evaluation results on the individual authorship style transfer task are shown in Table 1. We only show the Astrapop models trained with the full reward function $R=T+A-(LP^{\alpha}-1)$ since this reward function yields the best Joint score for all three PO algorithms on this task. The joint score indicates that the overall performance of the three Astrapop models are superior to all baseline models, and the two models trained with the RL-free PO algorithms (i.e. DPO and CPO) perform similarly to each other and both perform better than the RL-based PO algorithm (i.e. PPO).
Looking at the toward, away, and SBERT scores separately, we find that all PO algorithms can effectively improve the toward and the away scores, but at the cost of harming the SBERT score, since our reward function does not take semantic similarity into account, for efficiency and stability. Even so, the three Astrapop models still have decent SBERT scores that are higher than all baseline models except the SFT model since the KL-divergence penalty helps the models preserve the capability to keep the semantic meaning of the input texts. One may notice that STYLL has much better away and confusion scores than all other models, but this is because the model sometimes copies some irrelevant content from the target exemplars which changes the meaning of the transferred texts, and this also explains why the SBERT score for STYLL is much lower than other models.
@@ 4.2 Community Authorship Style Transfer
The automatic evaluation results on the community authorship style transfer task are shown in Table 2. For this task, the best Astrapop models are trained with the reward function without the away reward $R=T-(LP^{\alpha}-1)$. The joint score indicates that DPO- and CPO-Astrapop have the best overall performance. They also have the best toward and confusion scores. PPO can also slightly improve the performance of the SFT model, but the improvement is much less than DPO and CPO. Similar to the previous task, PO training harms the SBERT score, but the magnitude of the loss is very small, and the result SBERT scores are still higher than all baseline models except STRAP/SFT.
Method Toward Away SBERT Joint Confusion
Zero-shot 0.022 0.880 0.738 0.321 0.033
STYLL 0.210 0.832 0.854 0.598 0.227
STRAP / SFT 0.286 0.785 0.917 0.659 0.300
STEER 0.334 0.926 0.879 0.699 0.348
Astrapop-PPO 0.299 0.800 0.905 0.665 0.313
[…]
@@ 5 Related Work
Text Style Transfer Since parallel data is very rare for text style transfer, only a few works solve this task in a supervised manner (Zhu et al., 2010; Rao and Tetreault, 2018; Kim et al., 2022; Raheja et al., 2023). Constrained by the datasets, these works only focus on some specific sub-tasks such as text simplification and formality transfer. Therefore, to build more general style transfer models, recent works develop unsupervised methods that do not rely on parallel data. These works mainly fall in five categories, content-style representation disentanglement (Liu et al., 2019a; Jin et al., 2020), style-related phrase replacement (Madaan et al., 2020; Malmi et al., 2020; Reid and Zhong, 2021), reinforcement learning on direct objective (Gong et al., 2019; Liu et al., 2021b; Deng et al., 2022; Hallinan et al., 2023), pseudo-parallel data generation (Krishna et al., 2020; Riley et al., 2021), and LLM prompting (Reif et al., 2022; Suzgun et al., 2022; Patel et al., 2023).
The state-of-the-art authorship style transfer model, STYLL (Patel et al., 2023) transfers the input texts by prompting an LLM with the target style descriptors and few-shot pseudo-parallel transfer pairs generated by the same LLM, which combines the strength of pseudo-parallel generation and LLM prompting. Even so, as a prompting-based method, STYLL can be potentially enhanced by RL since RL has already been shown to be effective in improving the performance of prompting-based style transfer models (Deng et al., 2022), and the state-of-the-art general style transfer model, STEER (Hallinan et al., 2023) is also trained with RL. However, RL algorithms are shown to be unstable and hard to tune compared to the recently developed RL-free policy optimization algorithms such as DPO (Rafailov et al., 2023) and CPO (Xu et al., 2024).
Therefore, in this work, we choose the solve the authorship style transfer task with a PO-based training framework. Similar to STEER, we first generate pseudo-parallel data from the labeled non-parallel data and then train the model on the generated data, but our framework differs from STEER in three major ways: (1) we use a much simpler data generation strategy which only needs one paraphrase model and generates once for each instance in the non-parallel data, but STEER requires two extra models for each style as well as heavy overgeneration and filtering; (2) we only perform a single stage PO training instead of the two-stage offline-then-online RL training in STEER, and our reward function requires only one reward model instead of the three reward model in STEER; (3) we also use more stable and efficient RL-free PO algorithms instead of just the RL-based algorithm in STEER.
@@ 6 Conclusion
Although our approach shows strong performance on authorship style transfer, the performance on low-resource transfer is still much weaker than the performance on high-resource transfer. There are two possible reasons. First, we use small-scale datasets for both tasks due to the limited computational resources. It is sufficient to model the coarse-grained community authorship styles but may be insufficient for the individual authorship styles. Therefore, if more computational resources are available, future work can investigate whether more training data can help improve the performance of the low-resource authorship transfer models. Second, our authorship information injection strategy may not be optimal. We use a popular exemplar-based approach to inject the authorship information in the low-resource transfer task, but there may be more efficient approaches such as using continuous vectors instead of discrete tokens. This is out of the scope of this work, but future work can explore more efficient information injection strategies for low-resource authorship style transfer.
Moreover, even though the two RL-free PO algorithms, DPO and CPO already show a much better performance than PPO, in this work, we only use them in an offline manner as in the original papers. However, one can naturally enhance DPO and CPO with online data generated by the updated policy during training, which can potentially improve the performance of the models. Therefore, future work can focus on improving the training framework with online DPO or CPO training.
@@
Like other transfer learning LLMs, the quality of our model outputs highly depends on the quality of the underlying LLM and the training data. In this work, we use the original LLaMA-2-7B model instead of the chat version to ensure the flexibility for training, but it also has a higher risk of generating toxic texts. Also, the datasets we use contain unfiltered texts from the online forum Reddit and may also lead to unethical generation. Therefore, for real-world applications, we suggest carefully filtering the training data and also using a post-generation filter to avoid outputting unethical texts. As a PO-based training framework, one can also add some terms to the reward function to encourage the model to generate safe and ethical outputs.
Both datasets we use in this work contain texts with personal identifiable information (PII) and/or unethical words. We do not remove profane texts and texts containing PIIs for human evaluation to maximally preserve the style and meaning of the texts. Our human study protocol has been approved by an institutional review board.
Our model is intended for personal and authorized use such as building personal chatbots or authorship privatization, but we also recognize some potential harmful usage such as maliciously mimicking some individuals without authorization and intentionally generating texts in an offensive style. Therefore, we suggest keeping all personal data locally to prevent malicious mimicking. For text privatization, we suggest transferring to community-level authorship styles or styles mixed from multiple authors to prevent exposing the information of individual authors. To maximally preclude any unintended use, we only permit the use of our approach on public datasets or with the explicit consent of the target authors.